OpenStudy (anonymous):

What value of x makes (2x + 17) 1/3 = 3 a true statement?

OpenStudy (nikato):

1. Multiply both sides by 3. 2. Subtract 17 on both sides. 3.divide both sides by 2

OpenStudy (anonymous):

(2x + 17) 1/3 = 3 (2x+17)=9 2x=-8 x=-4

OpenStudy (anonymous):

THANKS!!!(: